Prove the following is a tautology (without using a truth table) [(p →q) (q + r)] → (p → r)
Answer #1

((p +9) ^ (q + r)) + (p + r) = -((pvq) ^ (Vr) v (p Vr) = -( p V 9) V-(-Vr) Vp Vs = (-1 9) ( -r) V-p V r. = pvp 9) ( -r) V r.